Find the value of k such that the graph of 4x = k -3y has a y- intercept of (0, -8)
Found 3 solutions by dabanfield, richwmiller, angphy:
Answer by dabanfield(803)

Find the value of k such that the graph of 4x = k -3y has a y- intercept of (0, -8)
Since (0,-8) is a point on the line we can substitute 0 for x and -8 for y in the equation:
4*0 = k - 3*-8
0 = k + 24
k = -24

Answer by richwmiller(17219)

0=k-3*-8
k=-24
4x=-24-3y
0=-24-3*(-8)

Answer by angphy(3)

4(0)=k-3(-8) since the y-intercept of (0,-8) means x=0 and y=-8, you substitute in those values
0=k-(-24)
0=k+24
-24 -24 subtract 24 from both side of the equation to isolate K (get the value of K)
k=-24